y=1/4x-7;(-8, ),( ,-5) i need the first and second ordered pair
please help someone

Answer by rwm(914)

I will assume you really mean y=(1/4)*x-7 or y=x/4-7
plug -8 in for x and solve for y
y=-8/4-7
y=-2-7
y=-9
(-8,-9)
plug -5 in for y and solve for x
-5=x/4-7
add 7
2=x/4
8=x
(8,-5)